OpenStudy (anonymous):

Hey guys.... I need help with this stupid problem. I don't know how to set it up. Here goes: You take a three question multiple choice test. Each question has four choices. You don't know any of the answers. What is the expirimental probability that you will guess exactly two out of three correctly?

OpenStudy (anonymous):

Chances that you guess one correctly = 1/4 Chances that you guess two correctly = 1/4 * 1/4 = 1/16 Chances that you guess two correctly and one false = 1/4 * 1/4 * 3/4 = 3/64

OpenStudy (anonymous):

oh that may not be right.

OpenStudy (anonymous):

Thanks, but the thing is it is EXPERIMENTAL probability. I got an online spinner. How would I set it up?

OpenStudy (anonymous):

2 correct and one wrong is not \[\frac{1}{4}\times \frac{1}{4}\times \frac{3}{4}\]

OpenStudy (anonymous):

because that is just the probability of (R, R, W)

OpenStudy (anonymous):

but you could also get (R, W, R) and also (W, R , R)

OpenStudy (anonymous):

so you have to multiply by 3, and get \[3\times \frac{1}{4}\times \frac{1}{4}\times \frac{3}{4}=\frac{9}{64}\]